Probability question*A box contains 3 white balls, 4 black balls, and 3 red balls. Consider selecting 3 balls at random. (a) What is the probability that you pick exactly one of each color when you select 3 balls from the box? (b) What is the probability that you pick exactly 2 white balls and 1 red ball? (c) What is the probability that at least one of the balls is white when you select 3 balls from the box?
Answer #1

a)

P(exactly one of each color)=inom{3}{1}inom{4}{1}inom{3}{1}/inom{10}{3} =3*4*3/120=36/120=3/10 =0.3

b)

P(exactly 2 whlte and 1 red ball)=inom{3}{2}inom{4}{0}inom{3}{1}/inom{10}{3}=3*1*3/120=9/120=3/40

c)

P(at least one white ball)=1-P(0 white ball)=00( 10 =1-1*35/120=1-7/24=17/24
Answer #2
? bar qmo me lele
source: Tere ma ke chod
answered by: Madarchod